Fred shovels snow twice as fast as Marvin. They require 16 minutes to clear the walk around their house when they work together. How long would it take Marvin to clear the walk working alone?
Answer by stanbon(75887)

Fred shovels snow twice as fast as Marvin. They require 16 minutes to clear the walk around their house when they work together. How long would it take Marvin to clear the walk working alone?
-----
Let Marvin time = 2x minutes/job ; rate = 1/(2x) job/min
----
Fred time = x minutes/job ; rate = 1/x job/min
---------
Together time = 16 min/job ; rate = 1/16 job/min
-----------------------
Equation:
rate + rate = together rate
1/2x + 1/x = 1/16
---
Multiply thru by 16x to get:
8 + 16 = x
x = 24 minutes (time needed by Fred)
2x = 48 minutes (time needed by Marvin)
==========================================
Cheers,
Stan H.